Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 114     Entry time: 27 May 2024, 06:41
Author: Konopkova Zuzana 
Group: HED 
Subject: Shift summary 

beam parameters: 2 colors at around 6.5 keV, 1 pulse

achievements:

user program continues with data acquisition at slightly varying photon energies (tens of eV both colors simultaneously) and at different delays between the colors

Issues:

  • users have scheduled datasets to be measured overnight with larger delays up to 350 fs, however, the BKR was able to go only up to 100 fs delay despite having larger delays earlier in the experiment (even during nights)
ELOG V3.1.4-7c3fd00